OpenStudy (anonymous):

One of the digits of a two-digit number is 3 less than the other digit. When the digits are interchanges, we get a new number. The sum of the squares of the new number and the initial number is 1877. Find the two numbers. I really don't know how to systematically go about doing this. I think a quadratic is involved though.

OpenStudy (kagıtucak):

let this numbers be 3x, x3 3x=30+x and x3=10x+3 then lets take square of this numbers (30+x)^2+(10x+3)^2=1877 900+60x+x^2+100x^2+60x+9=1877 101x^2+120x=968

OpenStudy (kagıtucak):

but i didnt find x , please check whether there is a mistake or not

OpenStudy (anonymous):

Thank you!

OpenStudy (nikato):

yes. he has made a mistake.. let one digit=a and the other=a+3

OpenStudy (nikato):

@simplethoughts so one number is 10(a+3)^2 + a the other is 10a+a+3

OpenStudy (nikato):

ignore the ^2 up there. so now [10(a+3)+a]^2 + [10a+a+3]^2= 1877 distribute (10a+30+a)^2 + (10a+a+3)^2 = 1877 combine like terms (11a+30)^2 +(11a+3)^2=1877 FOIL 121a^2+660a+900+121a^2+66a+9=1877 combine like terms and move everything to the left hand side 242a^2 +726a - 968 = 0 use quadratic formula and a=1 1+3=4, so other digit is 4 which means ur numbers are 14 and 41

OpenStudy (anonymous):

let number be xy either x = y-3 or x=y+3 (xy)^2 + (yx)^2 = 1877 Now it does not matter which one is bigger, because we are interchanging them tens place is gotten by mult by 10 [10(y-3)+y)]^2 + [10y+(y-3)]^2 = 1877 Perhaps this will work, but @nikato has solved it already.

OpenStudy (nikato):

@douglaswinslowcooper yup. got carried away. just got to continue and finish it with all my hard work

OpenStudy (anonymous):

Impressive solution.

OpenStudy (kagıtucak):

yeah i read the question wrongly but it is the logic anyway